OpenStudy (anonymous):

log(subscript 3)(5x-1)=log(subscript 3)(x+7)

OpenStudy (anonymous):

\[\log_{3}(5x-1)=\log_{3}(x+7) \]

OpenStudy (anonymous):

since they have the same bases just solve it as a normal equation 5x-1 = x+7 4x = 8 x=2

OpenStudy (jdoe0001):

$$\bf \large { log_{3}(5x-1)=log_{3}(x+7)\\ \text{using the cancellation rule of}\\ a^{log_ax} = x\\ \color{blue}{3}^{log_{\color{blue}{3}}(5x-1)}=\color{blue}{3}^{log_{\color{blue}{3}}(x+7)} \implies 5x-1 = x+7 } $$

OpenStudy (jdoe0001):

then as cmolina19 said, just solve for "x" :)

OpenStudy (anonymous):

You"re great teacher. I'm a new fan. Thank you for the help.